OpenStudy (anonymous):

Write the equation of the line passing through the points (0, -4) and (-2, 2).

OpenStudy (kamille):

Do you know formula of a line?

OpenStudy (anonymous):

no

OpenStudy (stamp):

Find the slope. Determine how y changes from one point to the next point, and do the same for x. So the slope is the ratio of the y change and the x change, or change in y / change in x Points (0, -4) and (2, -2). The y changes from -4 to -2, this means y changed + 2. The x changes from 0 to 2, this means x changed + 2. change in y / change in x = +2 / +2 the slope is +2 / +2, or just m = 1. Slope intercept form is y = mx + b where m is the slope. We know that m = 1. y = 1x + b y = x + b So what is b? Let us use a point to find it. (0, -4) is a point (x, y) on the line. Let us use that. y = x + b (-4) = (0) + b b = -4 y = x + b y = x - 4